Scottsdale Co. has actual sales for July and August and forecast sales for September, October, November, and December as follows:

Actual:
July $ 97,300
August 105,300
Forecast:
September 114,800
October 94,300
November 122,900
December 106,300

Based on past experience, it is estimated that 29% of a month’s sales are collected in the month of sale, 48% are collected in the month following the sale, and 9% are collected in the second month following the sale.

Required:

Calculate the estimated cash receipts for September, October, and November.

Cash collections
September
October
November
Answer #1

Answer

Cash collections
September $       92,593.00
October $       91,928.00
November $       91,237.00

Calculations:

September October November
29% of months sales $       33,292.00 $ 27,347.00 $ 35,641.00
48% of months sales $       50,544.00 $ 55,104.00 $ 45,264.00
9% of months sales $          8,757.00 $   9,477.00 $ 10,332.00
Total $       92,593.00 $ 91,928.00 $ 91,237.00

Formulae

September October November
29% of months sales =114800*29% =94300*29% =122900*29%
48% of months sales =105300*48% =114800*48% =94300*48%
9% of months sales =97300*9% =105300*9% =114800*9%

* It is assumed that remaining 14% (100- 29-48-9) are bad debts and are uncollectable, as there is no information given.

In case of any doubt, please comment.